### v1

# Enhancing the NAO: Extending Capabilities of Legacy Robots for Long-Term Research

Abstract
Many research groups face challenges when legacy (unsupported) robotic platforms lose manufacturer support and cannot accommodate modern sensing, speech, and interaction capabilities. We present the Enhanced NAO, a revitalized version of Aldebaran’s NAO robot that uses upgraded microphones, RGB-D and thermal cameras, and additional compute resources in a fully self-contained package. This system combines cloud and local models for perception and dialogue, while preserving the NAO’s expressive body and behaviors. In a pilot validation study, the Enhanced NAO delivered significantly higher conversational quality and stronger user preference compared to the NAO AI Edition, without increasing response latency. Key upgrades, such as beamforming microphones and low-latency audio processing, reduced artifacts like self-hearing and improved multi-party separation. Expanded visual and thermal sensing established a foundation for future interaction capabilities. Beyond the NAO, our framework provides a platform-agnostic strategy for extending the lifespan and research utility of legacy robots, ensuring they remain valuable tools for human-robot interaction.
I Introduction
The Aldebaran NAO [1] has been a staple in human–robot interaction (HRI) research for its approachable design, ease of use, and multimodal sensors. With Aldebaran’s recent bankruptcy and uncertainty around ongoing support, many labs risk losing a valuable research and education platform [2]. This challenge is not unique to NAO: Pepper was discontinued in 2021 [3], Rethink Robotics shut down in 2018, leaving Baxter obsolete [4, 5], and Willow Garage closed in 2014, ending support for the TurtleBot (first and second generations) and PR2 [6, 7, 8]. To address this broader problem of platforms becoming unsupported, we present a platform-agnostic framework for sustaining and enhancing obsolete robots, using the NAO as a central example.
Updating obsolete robots is challenging due to closed designs and outdated onboard compute resources. For NAO, prior work has added external computers [9], updated the operating system, [10], or integrated cloud services [11], but these approaches trade off portability, scope, or robustness.
We address these challenges with a fully integrated upgrade package (Fig. 1), featuring dual head-mounted cameras for multimodal perception, a Raspberry Pi Compute Module 5 (CM5) [12], a Seeed Studio ReSpeaker 4 Mic Array v2.0 [13] for robust speech pipelines, and a dedicated battery for peripheral power. This modular design extends sensing, processing, and interaction capabilities without reliance on manufacturer support or tethering.
Our contributions are threefold: (1) a complete hardware design, (2) an integration framework, and (3) supporting resources to be released open-source after review to enable widespread adoption. In a pilot validation study, the Enhanced NAO (with the custom upgrade package) demonstrated richer verbal interaction (e.g., conversation quality, participant preference) without added latency compared to the NAO AI, while expanded sensing and processing functionalities establish a practical blueprint for sustaining and enhancing legacy robots.
Figure 1: Enhanced NAO robot with integrated sensing, computing, and power systems. A. Back isometric view showing the ReSpeaker, custom CM5 carrier board, Hailo-8L, and portable battery pack. B. Front view highlighting the Orbbec Gemini 2 L depth camera, Seek thermal camera (S314SPX), and custom peripheral breakout board.
II Related Work
Researchers have extended the life of legacy research platforms (e.g., NAO [1], Baxter [4], PR2 [7], first and second generation TurtleBots [6]) through added compute resources [9, 14, 15], updated sensors [14, 16, 17, 18, 19, 20, 21], modified end-effectors [22, 17, 23], and alternative software stacks [10, 24, 25, 26, 27, 28, 29, 21].
For the NAO, researchers pursued software and hardware strategies[10, 24, 9, 14]. Software efforts replaced the proprietary NAOqi head-unit with Ubuntu [24] and ROS2 [10], but these require flashing new firmware[24] and are limited by the NAO’s small eMMC, requiring a permanently attached USB stick for operation [24]. Hardware approaches preserved the factory firmware but added external compute and peripherals via short, wired connections [9, 14] , enabling capabilities not supported by the stock head unit.
Similar modifications exist for other humanoid and non-humanoid robots. Prior software upgrades have replaced ROS1 with ROS2 for Baxter [25], abstracted low-level ROS2 interfaces for PR2 [27], and introduced new inverse kinematic libraries for smoother control of both robots [26, 28]. Hardware upgrades have included custom end-effectors [22, 17, 23, 30], clothing and tactile sensors to improve human-robot hugs [19], and RGB-D cameras for teleoperation and object detection [16, 17, 18]. Non-humanoid robotic platforms, such as the early TurtleBots have received comparable software [20, 29] and hardware upgrades [15, 21], extending their utility beyond their official support.
III Materials and Methods
Dimension 1 - Low 2 - Medium 3 - High
Relevance ($R$) Off topic or mostly unrelated Partially relevant but incomplete Fully relevant to the prompt and context
Specificity ($S$) Vague, generic Includes some surface level detail Clear and specific
Clarity ($C_{l}$) Hard to follow, confusing Understandable, some minor issues Concise, clear, and well-structured
Coherence ($C_{o}$) No logical flow from previous turn Mostly logical Flows naturally and connects to previous turns
Appropriateness ($A$) Flat tone Moderately engaging tone Warm, polite, and engaging tone
Timing ($t$) Long pause before response Robot interrupts user No issues
TABLE I: Rubric for evaluating conversation quality across six dimensions, each rated on a 1-3 scale, with higher scores indicating better quality.
Prior work demonstrates the potential and limitations of sustaining unsupported robots via piecemeal upgrades. These efforts highlight the need for integrated solutions. To address these limitations, we developed a platform-agnostic, custom hardware and software upgrade framework that introduces new data streams and enhances on-board compute resources for legacy, yet capable, robots. We demonstrate its usability on the NAO v6, a widely used platform in social and physical HRI research with uncertain long-term support.
III-A NAO Robot
We use the NAO v6 humanoid robot [1], a 58 cm-tall, 5.6 kg bipedal platform with 25 degrees of freedom (DOF), powered by an Intel Atom E3845 CPU with 4 GB RAM and a 32 GB eMMC storage running a Gentoo‑based Linux OS. The robot has two head-mounted HD cameras, a torso-mounted inertial measurement unit (IMU), and two ultrasonic sonars, four directional microphones, two speakers, nine capacitive touch sensors (three in the head, three per hand), and four bumper switches (two per foot). The platform supports Ethernet, Wi-Fi, and USB. Its lithium‑ion battery provides approximately 60-90 minutes of operation.
The NAO is controlled over a socket interface exposed by the proprietary NAOqi framework. Aldebaran provided a C++ library, Python bindings, and the Choregraphe GUI, though these require a local network connection to communicate with the socket interface. While the NAO v6 supports local code execution, it is restricted to Python 2.7, limiting developer flexibility [10].
RobotLab further markets the NAO AI Edition (hereafter referred to as “NAO AI”) with added cloud-based conversational features[31]. In contrast, our work builds on the base NAO v6 with custom hardware and software upgrades, collectively referred to as the “Enhanced NAO:”
III-B Hardware upgrades
III-B1 Computing and Power System
A Raspberry Pi CM5 serves as the main processor, featuring a Broadcom BCM2712 quad-core 64-bit Arm Cortex-A76 (Armv8) running at 2.4 GHz [12]. The CM5 is mounted on a carrier board with USB, a GPIO, and PCIe expansion, supporting a Hailo-8L accelerator for 13 TOPS of onboard vision-based inference [32].
Peripheral connections include USB ports and CSI/DSI interfaces for additional sensors (e.g., Seek Thermal Camera). The system is powered by a 5 V/5 A USB-C battery pack that sustains all peripherals. Mechanical enclosures, including backpack and sensor housings, were 3D printed, with mounts adapted from open-source designs [14] and custom components, as needed.ubsubsectionAudio System Audio capture is provided by the ReSpeaker with an embedded XMOS XVF-3000, offering built-in echo cancellation, noise suppression, and direction-of-arrival estimation [13].
III-B2 Camera System
The camera sensing suite integrates an Orbbec Gemini 2 L depth camera [33] and a Seek Thermal camera (S314SPX) [34]. The Orbbec provides synchronized RGB-D streams with native OpenCV compatibility, while the Seek Thermal delivers thermography data and RGB thermal visualizations.
Figure 2: Sample camera frames produced by the enhanced camera system. A) RGB and B) depth frames from the Orbbec Gemini 2 L and (C) thermal frame from the Seek Thermal (S314SPX) camera.
III-C Software upgrades
The CM5 runs Raspberry Pi OS with system nodes implemented in Python and C++.
III-C1 Camera Fusion
The CM5 processes RGB, depth, and thermal streams locally. Frames are synchronized and passed through the Hailo-8L for human detection, face region extraction, and landmark identification in thermal and RGB data. This pipeline establishes integration of thermal and visual cues for future interaction logic. Although not used in our pilot study, we implemented a demonstration pipeline integrating the Orbbec Gemini 2 L camera, Seek Thermal module, and the Hailo-8L (Fig. 2).
III-C2 Verbal Interaction Pipeline
Figure 3: Audio processing pipeline, starting with audio input via the ReSpeaker and ending with response audio played back through the NAO speakers.
The audio pipeline (Fig. 3) begins with the ReSpeaker, which outputs raw, processed, and echo-canceled streams. These signals are processed by the Open Embedded Audition System (ODAS) [35], for real-time speaker separation and sound source localization.
Deepgram [36] handles speech recognition, bypassing NAO’s built-in speech-to-intent engine to enable unconstrained dialogue. After transcription, LLM agents perform three steps: a) detecting whether the user is speaking, b) cleaning transcription errors, and c) identifying turn completion to refine the input before a final LLM agent generates the reply.
To reduce latency, responses are streamed immediately using Amazon Polly [37]. Because Polly does not natively support streaming, we developed a sentence-segmentation method that sends ordered requests, enabling a “think and speak” behavior and reducing the delay between user input and robot output.
III-D Pilot validation study
Figure 4: A participant engaging in a multi-turn conversation with the Enhanced NAO during the pilot validation study.
We conducted a within-subjects, counter-balanced pilot study with eight participants (four male presenting, four female presenting) to compare interactions with the NAO AI and the Enhanced NAO. Each 15-minute session included two conditions, with no compensation provided. In each condition, participants held a short conversation ($\geq 4$ turns; turn defined as a user utterance followed by a robot response). As shown in Fig. 4, the NAO was placed on a table facing the seated participant, bringing it closer to eye level. To ensure consistency, the first turn was always a user greeting to the robot; subsequent turns were open-ended, with the option to prompt the robot for a joke, if needed. After both conversations, participants indicated which interaction they preferred (e.g., in terms of ease and naturalness of conversation).
III-E Quantifying quality of responses
We assessed conversational quality using a two-level framework adapted from LLM evaluation methods [38]. The original framework included coherence, accuracy, clarity, relevance, and efficiency; we adapted it for conversational contexts by replacing accuracy with specificity, removing efficiency, and adding appropriateness and timing. This framework allows us to capture objective response quality and subjective user experience.
At the turn level, each robot response was rated on six dimensions (Table I): relevance (addresses user input), specificity (detail vs. vagueness), clarity (ease of understanding), coherence (logical flow), appropriateness (politeness and tone), and timing (delay before response).
At the conversation level, we computed mean scores for each dimension:
$\overline{D}=\frac{\sum_{i=1}^{n}d_{i}}{n}$ (1)
where $\overline{D}$ is the dimension mean, $d_{i}$ is the rating for the dimension at turn $i$, and $n$ is the number of turns rated. Once we obtain mean values for all six dimensions, we calculate an overall conversational quality score as:
$C_{q}=\frac{\overline{R}+\overline{S}+\overline{C_{l}}+\overline{C_{o}}+\overline{A}+\overline{t}}{6}$ (2)
where $C_{q}$ is the conversation quality score, and $\overline{R}$, $\overline{S}$, $\overline{C_{l}}$, $\overline{C_{o}}$, $\overline{A}$, and $\overline{t}$ represent the mean ratings for relevance, specificity, clarity, coherence, appropriateness, and timing, respectively.
We also examined turn balance, measuring whether conversations were dominated by the robot or the user. We define this ratio as:
$R=\frac{W_{r}}{W_{u}}$ (3)
where $W_{r}$ and $W_{u}$ denote the number of words spoken by the robot and the user, respectively. Their ratio quantifies relative talkativeness. Using Equation 3, an $R=1$ indicates balanced turn lengths, $R>1$ indicates robot-dominated speech, and $R<1$ indicates user-dominated speech.
Finally, we tracked topic shifts in each conversation, defined as major disruptions to conversational flow. A topic shift was recorded when the robot: 1) introduced an unrelated subject, 2) ignored or cut off the user’s input, stopping any elaboration, 3) returned to an earlier subject without linkage, or 4) closed a topic prematurely. While related to coherence, topic shifts capture disruptive changes in the semantic flow and direction of the conversation, whereas coherence addresses whether consecutive turns of the conversation flow logically. In this way, most topic shifts can be classified as coherence issues, but not all coherence issues are topic shifts.
IV Results
Feature NAO AI [31] Enhanced NAO[1] with Hardware/Software Upgrades
[…]
Frame rates may be impacted by NAO v6 compute resources.
In addition to the base NAO camera features:
[…]
Support for 2 additional CSI Cameras.
Microphones Four omnidirectional microphones; four channels at 48kHz or one channel at 16kHz. Accessible via NAOqi ALAudioDevice in 170ms chunks [1]. No algorithms to support ASR pipelines. Replaces existing NAO microphones
Seeed Studio ReSpeaker 4 Mic Array [13]
Six channels at 16kHz (4 raw, ASR, AEC)
Additional LED indicators.
Speech (TTS) Converts via NAOqi ALTextToSpeech. Supports pitch, speed, volume, pause, emphasis control, and voice effects. [39] Flexible with NAO’s default TTS engine, local TTS engines, or cloud-based TTS services. Tested implementation uses Amazon Polly [37].
Speech (ASR) Via NAOqi ALSpeechRecognition. Only recognizes predefined keywords. Requires a set vocabulary and returns matches with confidence scores. Supports word spotting within longer speech. [40] Flexible transcription using a dedicated ASR channel from the ReSpeaker. Tested implementation uses Deepgram to facilitate real-time speech-to-text.
Conversational
Capabilities
Static, intent-driven system focused on matching predefined intents with spoken keywords that trigger pre-determined responses. TTS engine has limited voices and lacks dynamic range. Supports open-ended dialogue with Deepgram [36], uses prompt engineering to tune an LLM [41] to generate relevant responses, and supports dynamic TTS voice generation via Amazon Polly [37].
TABLE II: Comparison of NAO AI capabilities and Enhanced NAO system features.
Figure 5: A comparison of average response times, conversation quality, turn balance, and number of topic shifts between the NAO AI (dark pink) and the Enhanced NAO (light pink). Boxes show the interquartile range (25th-75th percentiles), with the centerline representing the median and triangles marking the mean. Whiskers extend to the most extreme values not considered outliers. Circles indicate outliers. Black lines with asterisks indicate statistically significant differences after Bonferroni correction.
Figure 6: Participant preferences in the pilot study for interactions with the NAO AI (dark pink) versus the Enhanced NAO (light pink). Shading indicates participant gender: stripes indicate male, solid indicates female.
Table II summarizes the hardware and software improvements of the Enhanced NAO compared to the NAO AI.
IV-A Hardware
In addition to the NAO’s two head-mounted cameras, we integrated a depth and thermal camera, expanding multimodal vision with synchronized RGB-D and thermal streams. The Hailo-8L accelerator allows real-time inference with custom models on these streams, a capability unavailable on the NAO AI. The ReSpeaker handles audio capture, rather than the NAO v6’s built-in microphones, bypassing NAOqi’s restrictions. NAOqi provides audio only through ALAudioDevice in either 16 kHz or 48 kHz in 170 ms chunks [1], without support for beam steering or noise suppression, limiting custom ASR pipelines.
IV-B Software
We restructured the conversational pipeline by replacing the NAO’s intent-driven dialogue system with real-time, cloud-based transcription and diarization (Deepgram [36]). The Enhanced NAO uses an LLM, rather than fixed templates, for response generation, which enables dynamic, free-flowing conversation. Finally, to produce natural-sounding voices, we use Amazon Polly [37] for text-to-speech (TTS) generation.
IV-C Conversation
All conversations were coded by a primary researcher, with 20% independently coded by a secondary researcher to assess reliability. Inter-rater reliability [42], calculated using a two-way random intraclass correlation coefficient (ICC[2,1]), indicated excellent agreement between coders ($\text{ICC}=0.9897$).
For each of the four conversation attributes (average response time, conversational quality, turn balance, and number of topic shifts), we performed paired t-tests with a Bonferroni correction ($\alpha=0.0125$) to account for multiple comparisons (Fig. 5). We report effect sizes using Cohen’s $d$. Average response time did not differ significantly between the Enhanced NAO ($M=1.94,SD=0.42$) and the NAO AI ($M=2.13,SD=0.30$), $t(14)=1.05,p=0.3105,d=0.53$. Conversational quality was significantly higher for the Enhanced NAO ($M=2.89,SD=0.11$) compared to the NAO AI ($M=2.11,SD=0.37$), t(14)=5.69,p<0.001,d=2.85t(14)=-5.69,p<0.001,d=-2.85. Turn balance, was also significantly higher for the Enhanced NAO ($M=4.00,SD=1.19$) than the NAO AI ($M=0.86,SD=0.21$), t(14)=6.89,p<0.001,d=3.44t(14)=-6.89,p<0.001,d=-3.44, indicating that the Enhanced NAO produced substantially more words relative to the user. Additionally, the Enhanced NAO exhibited fewer topic shifts ($M=0.12,SD=0.35$) than the NAO AI ($M=1.44,SD=1.35$), $t(14)=2.66,p=0.0185,d=1.33$, which approached significance after the Bonferroni Correction.
Finally, user preference data (Fig. 6) shows that $87.5\%$ of users (seven out of eight) preferred the Enhanced NAO over the NAO AI, citing more natural and engaging conversation.
V Discussion
The hardware and software upgrades of the Enhanced NAO transformed it into a more capable and sustainable platform by adding sensing modalities, real-time inference, and low-level audio access. These enhancements move beyond the NAO AI’s capabilities and limitations to demonstrate how modern compute and perception capabilities can be layered onto legacy hardware. Although we did not benchmark the new camera and microphone system against the NAO’s integrated sensors, the additional RGB-D and thermal streams, together with advanced audio processing, provide valuable new data sources and greater developer control. Importantly, the upgrades were delivered in a fully self-contained, untethered package, demonstrating a platform-agnostic path for extending the lifespan and usability of unsupported robots.
These significant hardware upgrades directly improved the conversational quality of the Enhanced NAO. Cleaner, lower-latency audio from the ReSpeaker and the LLM-driven conversational pipeline generated more natural, context-aware responses. In contrast, the NAO AI’s rigid intent-based models produced repetitive, unnatural, and disruptive replies, reducing flow. Response times did not differ significantly between conditions, showing that richer interactions can be delivered without added latency, despite additional sensing, processing, and cloud services. Interestingly, turn balance skewed toward the Enhanced NAO, reflecting its ability to maintain a conversation, but also underscoring the need to preserve balance in future designs. Importantly, the Enhanced NAO showed fewer topic shifts, suggesting conversations stayed more coherent and user-driven.
V-A Limitations
This work was a pilot validation study to demonstrate the feasibility of our integrated system pipeline, rather than providing definitive evidence of the Enhanced NAO’s impact. The participant pool was necessarily small (eight participants), limiting the statistical power and generalizability of results. Nevertheless, we performed statistical analyses and reported effect sizes alongside significance tests, presenting results as transparently as possible and allowing readers to interpret the findings in context. Importantly, the large observed effects indicate clear promise and motivate follow-up studies with fully powered designs. For example, detecting a medium effect size in a within-subjects design comparing two robot conditions across two gendered populations would require approximately 36 participants, which we plan for future work.
A second limitation involved a technical inconsistency in one session, where a participant interacted with the Enhanced NAO using a male TTS voice instead of the female voice presented in all other sessions. While this mismatch may have influenced the system’s performance and the participant’s perception of the robot’s gender, it also revealed the flexibility of our pipeline to support multiple voices. Interestingly, this participant was the only one to prefer the NAO AI over the Enhanced NAO during qualitative feedback, which may have been due to the technical issue or the use of the male voice. Future experiments will ensure identical conditions across participants to eliminate such confounds.
Third, because this study emphasized integration over conversational design, aspects such as response prompt engineering and pretesting were not fully optimized, occasionally producing responses that were overly long or interview-like. These design refinements are straightforward to implement and will be addressed in subsequent studies, which will employ larger participant pools, standardized conditions, and improved dialogue strategies to rigorously evaluate system effectiveness.
Finally, we did not formally evaluate the new multimodal camera system in this pilot. Integration was the focus of this pilot, while assessing perceptual and interaction benefits will be pursued in future studies. These limitations reflect the preliminary nature of the study, but also underscore its success in validating the technical feasibility of our integrated framework and establishing the groundwork for more comprehensive evaluation.
VI Conclusion and future work
This work demonstrates a pathway for revitalizing legacy social robots by integrating modern sensing, speech, and dialogue capabilities, demonstrated with the NAO. Our system restored usability while delivering higher conversational quality and stronger user preferences, without added latency, compared to the NAO AI, highlighting the value of extending rather than retiring such platforms.
Future work will address the remaining technical challenges to further improve interaction quality. In particular, we will migrate our speech-to-text and text-generation pipelines from cloud services to local processing using compact, high-performance models such as Gemma 3n [43] or Qwen 3 [44]. Running these models locally is expected to further reduce latency, strengthen privacy, and increase system autonomy. We will also leverage multimodal visual perception to detect and localize users prior to speech, enabling natural dialogue initiation. In multi-party settings, fused visual and thermal data will support more accurate speaker separation by disambiguating overlapping voices and tracking active participants, further improving responsiveness.
Ultimately, we envision extending this framework beyond the NAO to demonstrate the platform-agnostic strategy for sustaining unsupported social robots. We are already deploying the framework to other legacy platforms like the Baxter and the TurtleBot 2, showcasing the system’s flexibility for humanoid and mobile robots. More broadly, this work highlights how thoughtful system integration can transform aging robots into modern research tools, extending their lifespan and ensuring they continue to advance human-robot interaction.
Acknowledgments
The authors thank the Ohio Space Grant Consortium for their generous support of the first author.

### v2

@@
 Abstract
-Many research groups face challenges when legacy (unsupported) robotic platforms lose manufacturer support and cannot accommodate modern sensing, speech, and interaction capabilities.
-We present the Enhanced NAO, a revitalized version of Aldebaran’s NAO robot that uses upgraded microphones, RGB-D and thermal cameras, and additional compute resources in a fully self-contained package.
-This system combines cloud and local models for perception and dialogue, while preserving the NAO’s expressive body and behaviors.
-In a pilot validation study, the Enhanced NAO delivered significantly higher conversational quality and stronger user preference compared to the NAO AI Edition, without increasing response latency.
-Key upgrades, such as beamforming microphones and low-latency audio processing, reduced artifacts like self-hearing and improved multi-party separation.
-Expanded visual and thermal sensing established a foundation for future interaction capabilities.
-Beyond the NAO, our framework provides a platform-agnostic strategy for extending the lifespan and research utility of legacy robots, ensuring they remain valuable tools for human-robot interaction.
+Legacy (unsupported) robotic platforms often lose research utility when manufacturer support ends, preventing integration of modern sensing, speech, and interaction capabilities.
+We present the Enhanced NAO, a revitalized version of Aldebaran’s NAO robot featuring upgraded beamforming microphones, RGB-D and thermal cameras, and additional compute resources in a fully self-contained package.
+This system combines cloud-based and local models for perception and dialogue, while preserving the NAO’s expressive body and behaviors.
+In a pilot user study validating conversational performance, the Enhanced NAO delivered significantly higher conversational quality and elicited stronger user preference compared to the NAO AI Edition, without increasing response latency.
+The added visual and thermal sensing modalities established a foundation for future perception-driven interaction.
+Beyond this implementation, our framework provides a platform-agnostic strategy for extending the lifespan and research utility of legacy robots, ensuring they remain valuable tools for human-robot interaction.
 I Introduction
-The Aldebaran NAO [1] has been a staple in human–robot interaction (HRI) research for its approachable design, ease of use, and multimodal sensors.
-With Aldebaran’s recent bankruptcy and uncertainty around ongoing support, many labs risk losing a valuable research and education platform [2].
-This challenge is not unique to NAO: Pepper was discontinued in 2021 [3], Rethink Robotics shut down in 2018, leaving Baxter obsolete [4, 5], and Willow Garage closed in 2014, ending support for the TurtleBot (first and second generations) and PR2 [6, 7, 8].
+The Aldebaran NAO [aldebaran_nao_2025] has been a staple in human–robot interaction (HRI) research for its approachable design, ease of use, and multimodal sensors.
+With Aldebaran’s recent bankruptcy and uncertainty around ongoing support, many labs risk losing a valuable research and education platform [strathearnSobolewska2025nao].
+This challenge is not unique to NAO: Pepper was discontinued in 2021 [nussey2021softbank], Rethink Robotics shut down in 2018, leaving Baxter obsolete [lawrenceCarol2019rethink, rethink2012baxter], and Willow Garage closed in 2014, ending support for the TurtleBot (first and second generations) and PR2 [garage2011turtlebots, garage2012pr2, willow_garage].
 To address this broader problem of platforms becoming unsupported, we present a platform-agnostic framework for sustaining and enhancing obsolete robots, using the NAO as a central example.
 Updating obsolete robots is challenging due to closed designs and outdated onboard compute resources.
-For NAO, prior work has added external computers [9], updated the operating system, [10], or integrated cloud services [11], but these approaches trade off portability, scope, or robustness.
-We address these challenges with a fully integrated upgrade package (Fig.
-1), featuring dual head-mounted cameras for multimodal perception, a Raspberry Pi Compute Module 5 (CM5) [12], a Seeed Studio ReSpeaker 4 Mic Array v2.0 [13] for robust speech pipelines, and a dedicated battery for peripheral power.
+For NAO, prior work has added external computers [mattamalaNaoBackpack2017], updated the operating system, [bono2024openaccessnaooan], or integrated cloud services [gestrin2024NaoChat], but these approaches trade off portability, scope, or robustness.
+We address these challenges with an integrated upgrade package (Fig.
+1), featuring dual head-mounted cameras for multimodal perception, a Raspberry Pi 5 [rpi5_2025], a Seeed Studio ReSpeaker 4 Mic Array [respeaker_2023] for robust speech pipelines, and a dedicated battery for peripheral power.
 This modular design extends sensing, processing, and interaction capabilities without reliance on manufacturer support or tethering.
-Our contributions are threefold: (1) a complete hardware design, (2) an integration framework, and (3) supporting resources to be released open-source after review to enable widespread adoption.
+Here, we contribute: (1) a complete hardware design, (2) an integration framework, and (3) supporting resources available upon request to enable widespread adoption.
 In a pilot validation study, the Enhanced NAO (with the custom upgrade package) demonstrated richer verbal interaction (e.g., conversation quality, participant preference) without added latency compared to the NAO AI, while expanded sensing and processing functionalities establish a practical blueprint for sustaining and enhancing legacy robots.
@@
 A.
-Back isometric view showing the ReSpeaker, custom CM5 carrier board, Hailo-8L, and portable battery pack.
+Back isometric view showing the ReSpeaker 4 Mic Array, Raspberry Pi 5, Hailo-8L, and battery pack.
 B.
-Front view highlighting the Orbbec Gemini 2 L depth camera, Seek thermal camera (S314SPX), and custom peripheral breakout board.
+Front view highlighting the Orbbec Gemini 2 L depth camera and Seek Thermal camera (S314SPX).
 II Related Work
-Researchers have extended the life of legacy research platforms (e.g., NAO [1], Baxter [4], PR2 [7], first and second generation TurtleBots [6]) through added compute resources [9, 14, 15], updated sensors [14, 16, 17, 18, 19, 20, 21], modified end-effectors [22, 17, 23], and alternative software stacks [10, 24, 25, 26, 27, 28, 29, 21].
-For the NAO, researchers pursued software and hardware strategies[10, 24, 9, 14].
-Software efforts replaced the proprietary NAOqi head-unit with Ubuntu [24] and ROS2 [10], but these require flashing new firmware[24] and are limited by the NAO’s small eMMC, requiring a permanently attached USB stick for operation [24].
-Hardware approaches preserved the factory firmware but added external compute and peripherals via short, wired connections [9, 14] , enabling capabilities not supported by the stock head unit.
+Researchers have extended the life of legacy research platforms (e.g., NAO [aldebaran_nao_2025], Baxter [lawrenceCarol2019rethink], PR2 [garage2012pr2], TurtleBots [garage2011turtlebots]) through added compute resources [mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet, bergeon2019raspberry], updated sensors [chatzilygeroudis2014naohelmet, avalos2017real, da2019control, paul2021object, block2019softness, song2019communication, macias2023map], modified end-effectors [devine2016real, da2019control, chu2015robotic], and alternative software stacks [bono2024openaccessnaooan, nao_image_2025, kermorgant2025baxterros2, kumar2024kinematics, wang2016pyride, ramezani2015smooth, alas2016controlling, macias2023map].
+For the NAO, researchers pursued software and hardware strategies[bono2024openaccessnaooan, nao_image_2025, mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet].
+Software efforts replaced the proprietary NAOqi head-unit with Ubuntu [nao_image_2025] and ROS2 [bono2024openaccessnaooan], but these require flashing new firmware[nao_image_2025] and are limited by the NAO’s small eMMC, requiring a permanently attached USB stick for operation [nao_image_2025].
+Hardware approaches preserved the factory firmware but added external compute and peripherals via short, wired connections [mattamalaNaoBackpack2017, chatzilygeroudis2014naohelmet] , enabling capabilities not supported by the stock head unit.
 Similar modifications exist for other humanoid and non-humanoid robots.
-Prior software upgrades have replaced ROS1 with ROS2 for Baxter [25], abstracted low-level ROS2 interfaces for PR2 [27], and introduced new inverse kinematic libraries for smoother control of both robots [26, 28].
-Hardware upgrades have included custom end-effectors [22, 17, 23, 30], clothing and tactile sensors to improve human-robot hugs [19], and RGB-D cameras for teleoperation and object detection [16, 17, 18].
-Non-humanoid robotic platforms, such as the early TurtleBots have received comparable software [20, 29] and hardware upgrades [15, 21], extending their utility beyond their official support.
+Prior software upgrades have replaced ROS1 with ROS2 for Baxter [kermorgant2025baxterros2], abstracted low-level ROS2 interfaces for PR2 [wang2016pyride], and introduced new inverse kinematic libraries for smoother control of both robots [kumar2024kinematics, ramezani2015smooth].
+Hardware upgrades have included custom end-effectors [devine2016real, da2019control, chu2015robotic, fitter2016equipping], clothing and tactile sensors to improve human-robot hugs [block2019softness], and RGB-D cameras for teleoperation and object detection [avalos2017real, da2019control, paul2021object].
+Non-humanoid robotic platforms, such as the early TurtleBots have received comparable software [song2019communication, alas2016controlling] and hardware upgrades [bergeon2019raspberry, macias2023map], extending their utility beyond their official support.
 III Materials and Methods
@@
 Coherence ($C_{o}$) No logical flow from previous turn Mostly logical Flows naturally and connects to previous turns
-Appropriateness ($A$) Flat tone Moderately engaging tone Warm, polite, and engaging tone
+Appropriateness (AA) Flat tone Moderately engaging tone Warm, polite, and engaging tone
 Timing ($t$) Long pause before response Robot interrupts user No issues
@@
 III-A NAO Robot
-We use the NAO v6 humanoid robot [1], a 58 cm-tall, 5.6 kg bipedal platform with 25 degrees of freedom (DOF), powered by an Intel Atom E3845 CPU with 4 GB RAM and a 32 GB eMMC storage running a Gentoo‑based Linux OS.
+We use the NAO v6 humanoid robot [aldebaran_nao_2025], a 58 cm-tall, 5.6 kg bipedal platform with 25 degrees of freedom (DOF), powered by an Intel Atom E3845 CPU with 4 GB RAM and a 32 GB eMMC storage running a Gentoo‑based Linux OS.
 The robot has two head-mounted HD cameras, a torso-mounted inertial measurement unit (IMU), and two ultrasonic sonars, four directional microphones, two speakers, nine capacitive touch sensors (three in the head, three per hand), and four bumper switches (two per foot).
@@
 Aldebaran provided a C++ library, Python bindings, and the Choregraphe GUI, though these require a local network connection to communicate with the socket interface.
-While the NAO v6 supports local code execution, it is restricted to Python 2.7, limiting developer flexibility [10].
-RobotLab further markets the NAO AI Edition (hereafter referred to as “NAO AI”) with added cloud-based conversational features[31].
+While the NAO v6 supports local code execution, it is restricted to Python 2.7, limiting developer flexibility [bono2024openaccessnaooan].
+RobotLab further markets the NAO AI Edition (hereafter referred to as “NAO AI”) with added cloud-based conversational features [robotlab_2022].
 In contrast, our work builds on the base NAO v6 with custom hardware and software upgrades, collectively referred to as the “Enhanced NAO:”
@@
 III-B1 Computing and Power System
-A Raspberry Pi CM5 serves as the main processor, featuring a Broadcom BCM2712 quad-core 64-bit Arm Cortex-A76 (Armv8) running at 2.4 GHz [12].
-The CM5 is mounted on a carrier board with USB, a GPIO, and PCIe expansion, supporting a Hailo-8L accelerator for 13 TOPS of onboard vision-based inference [32].
+A Raspberry Pi 5 serves as the main processor, featuring a Broadcom BCM2712 quad-core 64-bit Arm Cortex-A76 (Armv8) running at 2.4 GHz [rpi5_2025].
+The Raspberry Pi 5 provides USB, GPIO pins, and PCIe expansion, supporting a Hailo-8L accelerator for 13 TOPS of onboard vision-based inference [hailo20258l].
 Peripheral connections include USB ports and CSI/DSI interfaces for additional sensors (e.g., Seek Thermal Camera).
 The system is powered by a 5 V/5 A USB-C battery pack that sustains all peripherals.
-Mechanical enclosures, including backpack and sensor housings, were 3D printed, with mounts adapted from open-source designs [14] and custom components, as needed.ubsubsectionAudio System Audio capture is provided by the ReSpeaker with an embedded XMOS XVF-3000, offering built-in echo cancellation, noise suppression, and direction-of-arrival estimation [13].
-III-B2 Camera System
-The camera sensing suite integrates an Orbbec Gemini 2 L depth camera [33] and a Seek Thermal camera (S314SPX) [34].
+Mechanical enclosures, including backpack and sensor housings, were 3D printed, with mounts adapted from open-source designs [chatzilygeroudis2014naohelmet] and custom components, as needed.
+III-B2 Audio System
+Audio capture is provided by the ReSpeaker with an embedded XMOS XVF-3000, offering built-in echo cancellation, noise suppression, and direction-of-arrival estimation [respeaker_2023].
+III-B3 Camera System
+The camera sensing suite integrates an Orbbec Gemini 2 L depth camera [orbbec_gemini] and a Seek Thermal camera (S314SPX) [seek_2024].
 The Orbbec provides synchronized RGB-D streams with native OpenCV compatibility, while the Seek Thermal delivers thermography data and RGB thermal visualizations.
-Figure 2: Sample camera frames produced by the enhanced camera system. A) RGB and B) depth frames from the Orbbec Gemini 2 L and (C) thermal frame from the Seek Thermal (S314SPX) camera.
+Figure 2: Sample camera frames produced by the enhanced camera system.
+A) Processed RGB and B) depth frames from the Orbbec Gemini 2 L and (C) processed thermal frame from the Seek Thermal (S314SPX) camera.
 III-C Software upgrades
-The CM5 runs Raspberry Pi OS with system nodes implemented in Python and C++.
+The Raspberry Pi 5 runs Raspberry Pi OS with system nodes implemented in Node.JS, Python, and C++.
 III-C1 Camera Fusion
-The CM5 processes RGB, depth, and thermal streams locally.
+The Raspberry Pi 5 processes RGB, depth, and thermal streams locally.
 Frames are synchronized and passed through the Hailo-8L for human detection, face region extraction, and landmark identification in thermal and RGB data.
 This pipeline establishes integration of thermal and visual cues for future interaction logic.
-Although not used in our pilot study, we implemented a demonstration pipeline integrating the Orbbec Gemini 2 L camera, Seek Thermal module, and the Hailo-8L (Fig.
+Because the Enhanced NAO’s upgraded cameras augment rather than replace the existing NAO AI camera system, they were not used in the pilot study.
+Their integration expands sensing capabilities (adding depth and thermal perception) (Tab.
+II) without interfering with the existing system.
+However, we include a passive vision pipeline integrating the Orbbec Gemini 2 L camera, Seek Thermal module, and the Hailo-8L for demonstration (Fig.
 2).
 III-C2 Verbal Interaction Pipeline
-Figure 3: Audio processing pipeline, starting with audio input via the ReSpeaker and ending with response audio played back through the NAO speakers.
+Figure 3: Audio processing pipeline, from audio input via the ReSpeaker to a response audio played back through the NAO speakers.
 The audio pipeline (Fig.
 3) begins with the ReSpeaker, which outputs raw, processed, and echo-canceled streams.
-These signals are processed by the Open Embedded Audition System (ODAS) [35], for real-time speaker separation and sound source localization.
-Deepgram [36] handles speech recognition, bypassing NAO’s built-in speech-to-intent engine to enable unconstrained dialogue.
-After transcription, LLM agents perform three steps: a) detecting whether the user is speaking, b) cleaning transcription errors, and c) identifying turn completion to refine the input before a final LLM agent generates the reply.
-To reduce latency, responses are streamed immediately using Amazon Polly [37].
+These signals are processed by the Open Embedded Audition System (ODAS) [grondin2018lightweightoptimizedsoundsource], for real-time speaker separation and sound source localization.
+Deepgram [deepgram2025transcription] performs speech recognition and completion, bypassing NAO’s built-in speech-to-intent engine to enable unconstrained dialogue.
+Following Deepgram’s best practices, speech completion is detected when either the Endpointing feature identifies the end of natural speech in the raw audio stream or when the UtteranceEnd event indicates a sufficient silence gap in the transcript.
+Once a complete transcript is received, we align the word timings with ODAS outputs to produce speaker-aware transcripts.
+Although Deepgram achieves a low transcription error rate, performance can degrade with low speaker volume, microphone quality, distance, or background noise.
+Additionally, a complete transcript does not always indicate that a participant has finished speaking.
+To address these challenges, we created three GPT-4o [openai2024gpt]-based LLM agents with distinct roles: a) correcting transcription errors in speaker-aware transcripts, b) detecting turn completion, and c) generating a contextually appropriate response.
+We used prompt engineering to structure the system and user prompts for each agent, ensuring accurate task execution and coherent, contextually relevant replies that sustained natural, open-ended dialogue.
+To further reduce latency, responses are streamed immediately using Amazon Polly [AmazonPolly_2025].
 Because Polly does not natively support streaming, we developed a sentence-segmentation method that sends ordered requests, enabling a “think and speak” behavior and reducing the delay between user input and robot output.
 III-D Pilot validation study
-Figure 4: A participant engaging in a multi-turn conversation with the Enhanced NAO during the pilot validation study.
+Figure 4: A user engaging in a multi-turn conversation with the Enhanced NAO during the pilot validation study.
+Figure 5: Example of a participant conversation with the NAO AI and the Enhanced NAO.
 We conducted a within-subjects, counter-balanced pilot study with eight participants (four male presenting, four female presenting) to compare interactions with the NAO AI and the Enhanced NAO.
-Each 15-minute session included two conditions, with no compensation provided.
-In each condition, participants held a short conversation ($\geq 4$ turns; turn defined as a user utterance followed by a robot response).
+The study focused on evaluating conversational capabilities, the only functionality directly replaced and improved by the Enhanced NAO, using equivalent audio inputs and outputs.
+Each 15-minute session included two conditions, presented in counterbalanced order, with no compensation provided.
+In each condition, participants held a short conversation ($\geq 4$ turns; one turn defined as a user utterance followed by a robot response).
 As shown in Fig.
@@
 To ensure consistency, the first turn was always a user greeting to the robot; subsequent turns were open-ended, with the option to prompt the robot for a joke, if needed.
-After both conversations, participants indicated which interaction they preferred (e.g., in terms of ease and naturalness of conversation).
+A comparison of example conversations from both conditions is shown in Fig.
+5.
+After both conversations, each participant was verbally asked which conversation they preferred (e.g., in terms of ease and naturalness of conversation).
+Their responses were manually recoded by a research assistant.
+All sessions were video-recorded, manually transcribed, segmented into speaker turns, and timestamped for conversational quality assessment.
 III-E Quantifying quality of responses
-We assessed conversational quality using a two-level framework adapted from LLM evaluation methods [38].
+We assessed conversational quality using a two-level framework adapted from LLM evaluation methods [miller2025evaluating].
 The original framework included coherence, accuracy, clarity, relevance, and efficiency; we adapted it for conversational contexts by replacing accuracy with specificity, removing efficiency, and adding appropriateness and timing.
@@
 Their ratio quantifies relative talkativeness.
-Using Equation 3, an $R=1$ indicates balanced turn lengths, $R>1$ indicates robot-dominated speech, and $R<1$ indicates user-dominated speech.
+Using Equation 3, an $R=1$ indicates perfectly balanced turn lengths, $R>1$ indicates robot-dominated speech, and $R<1$ indicates user-dominated speech.
+This value is averaged across the entire interaction.
+Because natural conversation generally involves shared participation, we considered $0.5>R>2$ an acceptable range.
+Values outside this range signal issues in turn-taking, such as overly strict end-of-turn detection (user-dominated conversation) or overly verbose robot responses (robot-dominated conversation), both of which can lead to a less comfortable interaction.
 Finally, we tracked topic shifts in each conversation, defined as major disruptions to conversational flow.
@@
 In this way, most topic shifts can be classified as coherence issues, but not all coherence issues are topic shifts.
+Two evaluators (co-authors on this paper, who were not involved in the design or implementation of the Enhanced NAO) developed the coding schema before viewing any videos, to minimize expectation bias.
+All recordings were then coded by a primary researcher, with 20% independently coded by a secondary researcher to assess reliability.
+Although evaluators were aware of the condition assignments, all ratings followed standardized criteria to minimize bias.
 IV Results
-Feature NAO AI [31] Enhanced NAO[1] with Hardware/Software Upgrades
+Feature NAO AI [robotlab_2022] Enhanced NAO[aldebaran_nao_2025] with Hardware/Software Upgrades
 […]
@@
 Microphones Four omnidirectional microphones; four channels at 48kHz or one channel at 16kHz.
-Accessible via NAOqi ALAudioDevice in 170ms chunks [1].
+Accessible via NAOqi ALAudioDevice in 170ms chunks [aldebaran_nao_2025].
 No algorithms to support ASR pipelines.
 Replaces existing NAO microphones
-Seeed Studio ReSpeaker 4 Mic Array [13]
+Seeed Studio ReSpeaker 4 Mic Array [respeaker_2023]
 Six channels at 16kHz (4 raw, ASR, AEC)
@@
 Supports pitch, speed, volume, pause, emphasis control, and voice effects.
-[39] Flexible with NAO’s default TTS engine, local TTS engines, or cloud-based TTS services.
-Tested implementation uses Amazon Polly [37].
+[aldebaran_altexttospeech_2025] Flexible with NAO’s default TTS engine, local TTS engines, or cloud-based TTS services.
+Tested implementation uses Amazon Polly [AmazonPolly_2025].
 Speech (ASR) Via NAOqi ALSpeechRecognition.
@@
 Supports word spotting within longer speech.
-[40] Flexible transcription using a dedicated ASR channel from the ReSpeaker.
+[aldebaran_asr] Flexible transcription using a dedicated ASR channel from the ReSpeaker.
 Tested implementation uses Deepgram to facilitate real-time speech-to-text.
@@
 TTS engine has limited voices and lacks dynamic range.
-Supports open-ended dialogue with Deepgram [36], uses prompt engineering to tune an LLM [41] to generate relevant responses, and supports dynamic TTS voice generation via Amazon Polly [37].
+Supports open-ended dialogue with Deepgram [deepgram2025transcription], uses prompt engineering to tune an LLM (GPT-4o) [openai2024gpt] to generate relevant responses, and supports dynamic TTS voice generation via Amazon Polly [AmazonPolly_2025].
 TABLE II: Comparison of NAO AI capabilities and Enhanced NAO system features.
@@
 The ReSpeaker handles audio capture, rather than the NAO v6’s built-in microphones, bypassing NAOqi’s restrictions.
-NAOqi provides audio only through ALAudioDevice in either 16 kHz or 48 kHz in 170 ms chunks [1], without support for beam steering or noise suppression, limiting custom ASR pipelines.
+NAOqi provides audio only through ALAudioDevice in either 16 kHz or 48 kHz in 170 ms chunks [aldebaran_nao_2025], without support for beam steering or noise suppression, limiting custom ASR pipelines.
 IV-B Software
-We restructured the conversational pipeline by replacing the NAO’s intent-driven dialogue system with real-time, cloud-based transcription and diarization (Deepgram [36]).
+We restructured the conversational pipeline by replacing the NAO’s intent-driven dialogue system with real-time, cloud-based transcription and diarization (Deepgram [deepgram2025transcription]).
 The Enhanced NAO uses an LLM, rather than fixed templates, for response generation, which enables dynamic, free-flowing conversation.
-Finally, to produce natural-sounding voices, we use Amazon Polly [37] for text-to-speech (TTS) generation.
+To produce natural-sounding voices, we use Amazon Polly [AmazonPolly_2025] for text-to-speech (TTS) generation.
+These software modifications are part of the system’s functional improvements, which directly contribute to the Enhanced NAO’s observed gains in conversational quality and motivated the selection of our evaluation metrics.
 IV-C Conversation
-All conversations were coded by a primary researcher, with 20% independently coded by a secondary researcher to assess reliability.
-Inter-rater reliability [42], calculated using a two-way random intraclass correlation coefficient (ICC[2,1]), indicated excellent agreement between coders ($\text{ICC}=0.9897$).
-For each of the four conversation attributes (average response time, conversational quality, turn balance, and number of topic shifts), we performed paired t-tests with a Bonferroni correction ($\alpha=0.0125$) to account for multiple comparisons (Fig.
-5).
+As described in Sec.
+III, all conversations were independently coded by two researchers.
+Inter-rater reliability [hallgren2012computing], calculated using a two-way random intraclass correlation coefficient (ICC[2,1]), indicated excellent agreement between coders ($\text{ICC}=0.9897$).
+For each conversation attribute (average response time, conversational quality, turn balance, and number of topic shifts), we performed a paired t-test with a Bonferroni correction ($\alpha=0.0125$) to account for multiple comparisons (Fig.
+6).
 We report effect sizes using Cohen’s $d$.
@@
 Turn balance, was also significantly higher for the Enhanced NAO ($M=4.00,SD=1.19$) than the NAO AI ($M=0.86,SD=0.21$), t(14)=6.89,p<0.001,d=3.44t(14)=-6.89,p<0.001,d=-3.44, indicating that the Enhanced NAO produced substantially more words relative to the user.
-Additionally, the Enhanced NAO exhibited fewer topic shifts ($M=0.12,SD=0.35$) than the NAO AI ($M=1.44,SD=1.35$), $t(14)=2.66,p=0.0185,d=1.33$, which approached significance after the Bonferroni Correction.
+Additionally, the Enhanced NAO exhibited fewer topic shifts ($M=0.12,SD=0.35$) than the NAO AI ($M=1.44,SD=1.35$), $t(14)=2.66,p=0.0185,d=1.33$, which approached significance after the alpha correction.
 Finally, user preference data (Fig. 7) shows that $87.5\%$ of users (seven out of eight) preferred the Enhanced NAO over the NAO AI, citing more natural and engaging conversation.
@@
 Response times did not differ significantly between conditions, showing that richer interactions can be delivered without added latency, despite additional sensing, processing, and cloud services.
-Interestingly, turn balance skewed toward the Enhanced NAO, reflecting its ability to maintain a conversation, but also underscoring the need to preserve balance in future designs.
-Importantly, the Enhanced NAO showed fewer topic shifts, suggesting conversations stayed more coherent and user-driven.
+Interestingly, turn balance for the Enhanced NAO exceeded our desired range, yet participants still preferred it.
+This suggests that turn balance may be less critical to user preference than we initially assumed, or that our acceptable range was too conservative.
+Nevertheless, avoiding overly long or monologic robot responses is still necessary for maintaining enjoyable, user-centered conversation.
+Additionally, the Enhanced NAO exhibited fewer topic shifts, suggesting conversations stayed more coherent and user-driven.
+While the primary focus of this work was technical enhancement, we also acknowledge that the visible mechanical additions may have influenced user perception.
+The NAO’s humanoid, friendly appearance supports comfort and approachability, especially among children [shamsuddin2012humanoid, amirova202110, rossi2022using].
+Exposing additional hardware elements may make the robot appear more mechanical and potentially reduce approachability.
+Future adaptations should therefore consider visual and tactile design elements.
+Prior research on social and affective robots highlights the benefits of softness and warmth for comfort during social-physical interactions [block2019softness, block2017should, block2018emotionally, block2021six, block2023arms, block2021huggiebot].
+Inspired by this line of work, future iterations of the Enhanced NAO could incorporate a soft cover or hat over exposed components to preserve the NAO’s approachability while maintaining the improved functionality.
 V-A Limitations
@@
 A second limitation involved a technical inconsistency in one session, where a participant interacted with the Enhanced NAO using a male TTS voice instead of the female voice presented in all other sessions.
-While this mismatch may have influenced the system’s performance and the participant’s perception of the robot’s gender, it also revealed the flexibility of our pipeline to support multiple voices.
-Interestingly, this participant was the only one to prefer the NAO AI over the Enhanced NAO during qualitative feedback, which may have been due to the technical issue or the use of the male voice.
-Future experiments will ensure identical conditions across participants to eliminate such confounds.
-Third, because this study emphasized integration over conversational design, aspects such as response prompt engineering and pretesting were not fully optimized, occasionally producing responses that were overly long or interview-like.
-These design refinements are straightforward to implement and will be addressed in subsequent studies, which will employ larger participant pools, standardized conditions, and improved dialogue strategies to rigorously evaluate system effectiveness.
-Finally, we did not formally evaluate the new multimodal camera system in this pilot.
-Integration was the focus of this pilot, while assessing perceptual and interaction benefits will be pursued in future studies.
-These limitations reflect the preliminary nature of the study, but also underscore its success in validating the technical feasibility of our integrated framework and establishing the groundwork for more comprehensive evaluation.
+This mismatch may have influenced the system’s performance and the participant’s perception of the robot’s gender, and notably, this was the only participant who preferred the NAO AI.
+The deviation could also have interacted with turn-balance patterns.
+However, it also revealed the flexibility of our pipeline to support multiple voices.
+We did not analyze male vs.
+female voice as an experimental factor in this pilot, but future work should explicitly test how voice characteristics influence conversational quality and user preference.
+Future experiments will also implement stricter quality control to ensure consistent conditions across participants.
+Additionally, participants were not blinded to the experimental condition.
+To mitigate bias, we used a counter-balanced condition order and asked participants to evaluate the conversation quality rather than the robot itself.
+Future studies could incorporate blinded or partially blinded setups to minimize potential bias due to visual differences between conditions.
+Forth, because this pilot emphasized system integration over conversational design, prompt engineering and pretesting were not fully optimized, occasionally producing overly long or interview-like responses.
+These refinements are straightforward to implement and will be incorporated in future studies with larger participant pools, standardized conditions, and improved dialogue strategies.
+Finally, although the new multimodal camera system was successfully integrated, we did not formally evaluate its perceptual or interaction benefits in this pilot.
+Integration was the primary goal; follow-up work will characterize the contributions of these sensing modalities in depth.
+Taken together, these limitations reflect the preliminary nature of the study, but also underscore its success in validating the technical feasibility of our integrated framework and establishing the groundwork for more comprehensive evaluation.
 VI Conclusion and future work
@@
 Future work will address the remaining technical challenges to further improve interaction quality.
-In particular, we will migrate our speech-to-text and text-generation pipelines from cloud services to local processing using compact, high-performance models such as Gemma 3n [43] or Qwen 3 [44].
+In particular, we will migrate our speech-to-text and text-generation pipelines from cloud services to local processing using compact, high-performance models such as Gemma 3n [gemma_3n_2025] or Qwen 3 [qwen3technicalreport].
 Running these models locally is expected to further reduce latency, strengthen privacy, and increase system autonomy.